unseen:

Solve for x. 6x + 2 = 14 x = 1 x = 2 x = 2 x = 12

unseen:

heelp

iosangel:

subtract 2 from BOTH sides because we're doing the opposite of addition: 6x+2-2=14-2=6x=12 now divide "6" by BOTH sides like so: 6x/6=12/6= (the"6x/6" will be crossed out) and 12/6=2 Answer: x=2